Using the data provided below, please provide:

1. Contribution Margin (CM)

2. Contribution Margin Ratio (CMR)

3. Breakeven in Sales Dollars

4. Margin of Safety (MOS)

5. MOS ratio

    

Units sold:          20,000

Selling price: $             20

Variable manufacturing costs: $    140,000

Fixed manufacturing costs: $      92,000

Variable selling costs: $      50,000

Fixed administrative costs: $      15,000
Answer #1

--Requirements

A Sales revenue [20000 x $ 20] $400,000
B Variable manufacturing cost $140,000
C Variable administrative cost $15,000
D = A - B - C Contribution margin (CM) $245,000 Answer
E = (D/A) x 100 CM Ratio 61.25% Answer
F Fixed Cost $107,000
G = F/E Breakeven in Sales Dollars $174,694 Answer
H = A - G Margin of Safety Dollar $225,306 Answer
I = (H/A) x 100 MOS Ratio 56.33% Answer